OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8 (3, 2) (2, 3) (-2, -3) (-2, 3)

OpenStudy (anonymous):

x=-2y+8 4(-2y+8)+y=11 -8y+32=11 -8y=-20 y=2.5

OpenStudy (anonymous):

I made a mistake, -8y+32+y=11 -7y=-21 y=3

OpenStudy (anonymous):

Now that we have y, x will be easy to find

OpenStudy (anonymous):

Thanks i got it, its 2,3 :)!

OpenStudy (anonymous):

Yay :D